4. Ray Tucker, owner of Tucker Grocery, does all of the buying for his company. Recently, Coca-Cola offered Ray a deal on 2 pallets of Coke and Diet Coke for Homecoming weekend. Each pallet contains 50 24-pack cases (½ Coke, ½ Diet Coke.) Normally the cost per case is $4.00 and the retail is $5.00. Coca-Cola is now offering Ray a cost per case of $3.85.

Part 1 - If Ray would like to maintain the same initial markup % (as he had at the $4.00 cost) on the cases of Coke and Diet Coke, what price should he now sell the cases of Coke and Diet Coke for?

Part 2 - If Ray decided that he wanted to maintain the same “penny profit,” would his price for the cases of Coke and Diet Coke be the same as the amount you just calculated?
Part 3 - What will Ray’s dollar profit be on the cases of Coke and Diet Coke if he sells all the cases at the price you calculated in Part 1?
Answer #1

Solution:

Initially, with cost of $4 and retail price of $5 per case,

Markup = (price/cost) - 1 = 5/4 - 1 = 1.25 - 1 = 0.25 or 25%

Penny/cash profit = retail price - cost = $5 - $4 = $1 per case

1) Now, the cost = $3.85. For markup = 0.25, we need to find the retail price

So, 0.25 = (price/3.85) - 1

Or (0.25 + 1)*3.85 = price = $4.8125

So, to maintain the same initial markup, now the cases should be sold for $4.8125

2) We found penny profit to be $1. Had the penny profit was to be same, retail price should have been $3.85 + $1 = $4.85 (which is not equal to $4.8125). So, no, the price calculated above is lower than the price required to maintain same penny profit.

3) We know that each case is sold for $4.8125, and there are 50 such cases in a pallet and two such pallets. So, total number of cases = 50*2 = 100 cases

Profit = total revenue - total cost

total revenue = price per case*number of cases and total cost = cost per case*number of cases

Profit = 100*4.1825 - 100*3.85 = $96.25